# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v2

@@ III.1 Baseline Flow
Fig. 5 depicts the vortex shedding patterns within the wake field of elliptical cylinders with $Ar$ from 0 to 2. Starting from Fig. 5(a) depicting an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding. This is expected as the elongated elliptical cylinder results in weak adverse pressure gradients that allow the boundary layers to stay attached for a prolonged distance. The zero vorticity contour lines in the wake extend all the way to the outlet with slight fluctuations. This suggests that the wake is largely stable, with minor flow instability due to shear in the flow. It is worth highlighting that this test case of $Ar=2$ serves as a special verification study for subsequent DRL control experiments. For a stable flow field with no vortex shedding, we expect the DRL-based flow control algorithm to conform to the underlying flow physics and to not introduce any actuation since the current baseline is essentially the optimal flow configuration. To the best of our knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in literature. We include this to showcase the robustness of DRL-based control across a spectrum of flow physics. We delay the discussion of DRL control to Sec. III.2.
Figure 5: Contours of instantaneous vorticity for flow around an elliptic cylinder at $Re=100$, varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
@@
Velocity fluctuation and design of probe layout
In addition to the baseline flow field, we analyze the fluctuating component of velocity, which serves as a good indicator to inform the location of probes during control. The judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al. (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes. We also utilize the physical information from the wake field to guide the flow control strategies of the agents. This aspect is particularly reflected in the practical design of the probe positions. The study by Protas and Wesfreid Protas and Wesfreid (2002) elaborates on the adjustable component of the drag coefficient, $C_{D}^{0}$, which reflects the influence of oscillatory flow phenomena on the drag coefficient. This provides a pathway for strategic manipulation of the von Kármán vortex street to adjust the drag through targeted interventions. Building on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=0.25$; (b) $Ar=0.1$; (c) $Ar=0$.
As depicted in Fig. 6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0, 0.1, and 0.25 are illustrated. The positions of the black dots on the diagram represent the locations of the probes. These probes are strategically placed around the elliptical cylinder and in the downstream wake region. On one hand, the arrangement of probes around the elliptical cylinder is designed to capture physical information pertinent to the lift and drag forces, which are critical components of our control objectives. On the other hand, probes within the wake of the elliptical cylinder are positioned to encompass areas of maximum velocity fluctuations, facilitating the extraction of flow field data from regions where velocity fluctuations and vortex instabilities are most intense. The extraction of physical information from these locations enables the agents to acutely observe changes in the flow field. In Fig. 6, three tests with different $Ar$ are shown as examples. The configurations for other $Ar$ are fundamentally similar and are not elaborated further.
@@ III.2 Effectiveness of AFC
This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control. At $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, from elongated body to circular cylinder and then to a normal flat plate. The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process. Fig. 7 and Fig. 8 display the velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0 under baseline and controlled conditions. More specifically, Fig. 7 shows those for $Ar$ from 2 to 0.75, the vortices in the wake field of the controlled elliptical cylinders can be completely suppressed. In contrast, Fig. 8 illustrates those for $Ar$ from 0.5 to 0, the vortices in the wake field of the controlled elliptical cylinders are mitigated by not fully suppressed.
@@
Starting from Fig. 7(a) for the elliptical cylinder with $Ar=2$, the baseline flow field (left) exhibits high stability with no observable vortex shedding. We applied a DRL-based flow control to this stable, non-shedding wake field, and the controlled flow field (right) behind the elliptical cylinder remained stable and nearly identical to the baseline. This sanity test indicates that the DRL-based model is consistent with flow physics and does not generate any unreasonable perturbation during the control process. As the $Ar$ decreases to $Ar=1.75$, the application of DRL-based AFC technology successfully suppresses the vortices that originally separated alternately from both sides of the elliptical cylinder. Further reducing the $Ar$ results in a more pronounced and frequent shedding of vortex structures in the uncontrolled wake, while the recirculation bubble downstream of the elliptical cylinder gradually decreases and becomes oscillatory. Although the flow instability gradually increases, under the conditions of $Ar=1.5$, 1.25, 1, and 0.75, the alternating vortex streets behind the elliptical cylinders are transformed into a stable wake field after the implementation of AFC in all cases. The DRL-based AFC technology demonstrates significant vortex shedding suppression capabilities under these conditions, highlighting the effectiveness of the AFC control strategy in enhancing the flow stability of the wake field.
@@
In Fig. 8, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the baseline flow becomes increasingly unstable, and the area of the recirculation region gradually decreases while the strength and frequency of vortex shedding increase. When $Ar=0.5$, the frequent alternating vortex shedding is effectively suppressed after flow control is applied. The recirculation region behind the elliptical cylinder is significantly elongated, and the shedding of vortices is controlled, although the wake flow still exhibits alternating oscillations. For the controlled flow at $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is elongated, resembling the case at $Ar=0.5$. However, there is a notable difference: the recirculation bubble now demonstrates prominent oscillations. The vortex shedding behind the cylinder continues alternately, akin to the elongated vortex structures being stretched longitudinally. Despite these instabilities, vortex shedding in the wake fields at $Ar=0.5$ and $Ar=0.25$ is reasonably mitigated compared to the baseline state. At $Ar=0.1$ or 0, compared to baseline conditions, the recirculation bubbles behind the cylinder are elongated, and the initiation point of vortex shedding is extended further downstream. We do not observe complete suppression of vortex shedding due to the highly chaotic nature of these wake behind vertical thin bodies. Nevertheless, the vortex shedding is reduced to a lower frequency and intensity. These observations reflect the effectiveness of the control strategies in suppressing vortex shedding and stabilizing the wake field.
@@
Fig. 9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75. At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum. After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value. Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab. 2. These observations substantiate the effectiveness of the AFC strategy underpinned by DRL algorithms. The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages. This demonstrates the capability of the DRL-based AFC approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
Fig. 10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0. Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher. Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes. However, they did not fully stabilize in the subsequent period. In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow. At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow. Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced. Similarly, the $C_{L}$ undergoes an initial sharp decrease and approaches near zero, then later starts to oscillates, but at a lower frequency than the baseline $C_{L}$ oscillation.
@@
Figure 10: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color indicates that the drag or lift coefficient is still in an unstable stage after flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
Figure 11: Action, which is the mass flow rate of the synthetic jets. The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
Fig. 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2. When $Ar=2$, the AFC is negligible, and the value of the action, pertaining to the control input, remains close to zero. This is a strong indication that the DRL-based model is not generating nonphysical perturbations for a stable flow field. From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets. This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs. As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets. Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets. This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input. When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages. For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
Overall, the results presented in Fig. 7 and Fig. 9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=2$ to $Ar=0.75$. By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input. However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes. At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab. 2.
In Fig. 12, we observe the learning curve during training of flow around elliptical cylinders with different $Ar$. At $Ar=2$, the learning curve has entered the plateau phase after the initial improvement phase, indicating that the DRL training has converged. For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4,000 episodes converges near the maximum reward value. At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3,000 episodes and eventually settles into a stable plateau. For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning. After 2,000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2,000 and 5,000 episodes. This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance. At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4,000 episodes, with the reward values stabilizing between 4,000 and 10,000 episodes. At $Ar=0.25$, training approaches convergence after 8,000 episodes, reaching a total of 20,000 episodes, with a mid-to-late plateau marking stable convergence. At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12,000 episodes.
The DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process. Training also achieves stable convergence within 5,000 episodes for $Ar=1.5$ and $Ar=1.75$. However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed. This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
@@
Table 2: The control effect of elliptical cylinder under DRL-based AFC.
@@
  • a. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged drag coefficient of the elliptical cylinder once the baseline flow has fully developed.
  • b. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase.
  • c. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
Tab. 2 presents the effectiveness of AFC techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0 to 2. It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change. Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate. When the $Ar$ is set to 2, the baseline flow has reached a steady state. During the application of flow control, the agent does not use external energy to disrupt this stable flow field, which demonstrates the intelligence of the DRL-based AFC. This reflects the adaptability and intelligent decision-making of AFC control in maintaining flow stability. For elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case. These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy. However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively. This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder. As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process. The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig. 8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity. These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
[…]
@@
In their research, Protas and WesfreidProtas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations. This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control. In our research, the flow field is controlled by utilizing external excitation, such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag. This is the reason why the drag reduction rate gradually increases as $Ar$ decreases in Tab. 2.

### v3

@@ III.1 Baseline Flow
Fig. 5 depicts the vortex shedding patterns within the wake field of elliptical cylinders with $Ar$ from 0 to 2. For an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding. This is expected as the elongated elliptical cylinder results in weak adverse pressure gradients that allow the boundary layers to stay attached for a prolonged distance. The zero vorticity contour lines in the wake extend all the way to the outlet with slight fluctuations. This suggests that the wake is largely stable, with minor flow instability due to shear in the flow. It is worth highlighting that this test case of $Ar=2$ serves as a special verification study for subsequent DRL control experiments. For a stable flow field with no vortex shedding, we expect the DRL-based flow control algorithm to conform to the underlying flow physics and to not introduce any actuation since the current baseline is essentially the optimal flow configuration. Based on our current knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in the literature. We include it to emphasize the robustness of DRL-based control across a spectrum of flow physics. The detailed discussion of DRL control will be presented in Sec. III.2, where we will provide a comprehensive analysis of its principles and application.
Figure 5: Contours of instantaneous vorticity for flow around an elliptic cylinder at $Re=100$, varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. Purple represents the clockwise vortices shed from the upper side of the cylinder, while orange represents the counterclockwise vortices shed from the lower side. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1.0$; (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (g) $Ar=0$.
@@
Velocity fluctuation and probe layout optimization
In addition to investigating the baseline flow field, we conduct an analysis of the velocity fluctuation component, which serves as a valuable guiding indicator for probe placement during control. By integrating our knowledge of fluid physics into the control agents, we can optimize their architecture while maintaining superior performance. An illustrative example of this approach is demonstrated in the work of Pastoor et al.Pastoor et al. (2008), where phase control for drag reduction of a bluff body is achieved by using the decoupled development of shear layers and wake processes. Similarly, we utilize the physical information from the wake field to guide the flow control strategies of the agents, with a focus on the practical design of probe positions. Moreover, Protas and Wesfreid Protas and Wesfreid (2002) highlighted the influence of oscillatory flow phenomena on the drag coefficient, represented by the adjustable component $C_{D}^{0}$. This understanding provides a strategic pathway for manipulating the von Kármán vortex street and achieving drag adjustment through targeted interventions. Expanding on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=1$; (b) $Ar=0$. To provide a clearer visualization of the probe distribution, we display to a specific region within the computational domain. The x-axis ranges from -2D to 10D, and the y-axis ranges from -2D to 2.1D, where D represents a reference length scale.
As depicted in Fig. 6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0 and 1 are illustrated. The positions of the black dots on the diagram represent the locations of the probes. These probes are strategically placed around the elliptical cylinder and in the downstream wake region. On one hand, the arrangement of probes around the elliptical cylinder is designed to capture physical information pertinent to the lift and drag forces, which are critical components of our control objectives. On the other hand, probes within the wake of the elliptical cylinder are positioned to encompass areas of maximum velocity fluctuations, facilitating the extraction of flow field data from regions where velocity fluctuations and vortex instabilities are most intense. The extraction of physical information from these locations enables the agents to acutely observe changes in the flow field. In addition, Fig. 6 illustrates two tests with different $Ar$, the configurations for other $Ar$ are similar and not further elaborated upon.
@@ III.2 Effectiveness of AFC
Using the DRL algorithm, we investigate the performance of actively controlling the magnitude of synthetic jets for targeted flow control throughout the process of elliptical cylinders transitioning from elongated bodies to circular cylinders, and finally to a normal flat plate, with aspect ratios ranging from 2 to 0 at $Re=100$. The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in $C_{D}$, weakening of $C_{L}$, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
For a detailed demonstration of vortex suppression, we present in Fig. 7 an example using an elliptical cylinder with $Ar=0.75$, observing the evolution of the instantaneous vorticity field throughout the DRL training process. We select four representative snapshots of the instantaneous vorticity field to depict changes in vortex shapes and positions. Labels A,B,C,D and E in Fig. 7(a) denote the five vortices initially shed. By the time depicted in Fig. 7(b), these initial five vortices have progressed downstream without new vortices being shed under the influence of flow control. In Fig. 7(c), labeled as A,B,C,D and E, the vortices have developed near the exit of the computational domain, exhibiting elongated shapes as they are about to be expelled from the domain. At this stage, the shape of the recirculation bubble has stabilized, and no new vortices are shed. By the time represented in Fig. 7(d), the originally labeled five vortices have all been expelled from the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain has reached a stable state. Under the conditions of $Ar=1.75$, 1.5, 1.25, and 1 for the elliptical cylinder, the DRL-based AFC demonstrates similar control effectiveness.
Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training. Specifically, focused on four representative training episodes: (a) Episode = 0, (b) Episode = 450, (c) Episode = 800, and (d) Episode = 2100. The initial shedding vortex are labeled as A, B, C, D and E, and them deformation and displacement are subsequently tracked. The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
@@
To demonstrate the suppression effect of AFC technology on vortex shedding phenomena, velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0.75 and from 0.5 to 0 are displayed in Fig. 8 and Fig. 9, respectively. As shown in Fig. 8, the DRL-based AFC approach demonstrates consistent and effective suppression of vortex shedding and enhancement of flow stability in the wake field. At $Ar$ = 2, the baseline flow field is stable without vortex shedding, and the controlled flow field remains stable and similar to the baseline. As $Ar$ decreases, the AFC successfully suppresses vortices that were previously formed on both sides of the elliptical cylinder. Despite increased vortex shedding and oscillatory behavior in the uncontrolled wake, AFC implementation at $Ar$ values of 1.5, 1.25, 1, and 0.75 transforms the alternating vortex streets into a stable wake field. Subsequently, as the $Ar$ continuously decreases, the baseline flow becomes more unstable, characterized by a diminishing recirculation region and increased vortex shedding intensity and frequency. This phenomenon is visually depicted in Fig. 9.
In particular, at $Ar$ = 0.5, the application of flow control successfully suppresses the alternating vortex shedding, resulting in an elongated recirculation region. However, the wake flow still exhibits oscillations. For $Ar$ = 0.25, flow control elongates the recirculation bubble similar to the $Ar$ = 0.5 case, but with prominent oscillations. Vortex shedding persists in an alternating manner, resembling elongated and stretched vortex structures. Although complete suppression of vortex shedding is not achieved at $Ar$ = 0.1 or 0 due to the inherent chaotic nature of the wake behind vertical thin bodies, the recirculation bubbles are elongated, and the initiation point of shedding is shifted further downstream. The controlled flow exhibits reduced shedding frequency and intensity compared to the baseline conditions. Fig. 8 illustrates the effectiveness of AFC technology in eliminating vortex shedding and transitioning flow from unstable to stable modes within the range of $Ar$ from 2 to 0.75, while Fig. 9 indicates that at lower $Ar$ values (0.5 to 0), this technology can effectively alleviate vortex shedding and enhance flow stability to some extent.
@@
Further, Fig. 10 and Fig. 11 provide a comparative analysis of the $C_{D}$ and $C_{L}$ before and after implementing the DRL-based AFC strategy, showcasing its effectiveness across a range of $Ar$ from 2 to 0.75 and from 0.5 to 0, respectively. As shown in Fig. 10, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2. However, when synthetic jets control is activated at $Ar$ values of 1.75, 1.5, 1.25, 1, and 0.75, there is an immediate and substantial reduction in $C_{D}$, which quickly reaches a minimum value. After half of the total control duration, $C_{D}$ stabilizes and remains constant at this minimum value. Similarly, $C_{L}$ experiences significant fluctuations initially but stabilizes near zero after half of the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ values can be found in Tab. 2. Clearly, the periodic oscillations of $C_{D}$ and $C_{L}$ originally present in the baseline flow are significantly reduced, ultimately transitioning to a stable state.
@@
Figure 11: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color indicates that the drag or lift coefficient is still in an unstable stage after flow control. (a) and (b) $Ar=0.5$; (c) and (d) $Ar=0.25$; (e) and (f) $Ar=0.1$; (g) and (h) $Ar=0$.
The impact of the AFC strategy on $C_{D}$ and $C_{L}$ at $Ar=0.5$ and $Ar=0.25$ is described in Fig. 11. While both coefficients experience reductions compared to the baseline flow, they do not achieve the same level of stability observed at $Ar=0.75$ and higher. Initially, there are abrupt changes in both $C_{D}$ and $C_{L}$, but they do not fully stabilize in the subsequent period. In the latter half of the control duration, there are fluctuations in both coefficients, but with lower amplitude and frequency compared to the baseline flow. At $Ar=0.1$ and $Ar=0$ are described in also in Fig. 11, the controlled flow exhibits a sharp decrease in drag coefficient at the beginning of the control, followed by oscillations at a lower frequency and amplitude compared to the baseline flow. Although complete stability is not achieved, the average value of $C_{D}$ is significantly reduced. Similarly, $C_{L}$ initially decreases sharply and approaches zero, then exhibits oscillations at a lower frequency than the baseline $C_{L}$ oscillations.
@@
Table 2: Quantified control effects of elliptical cylinders under DRL-based AFC.
@@
The DRL-based AFC implemented in our study is demonstrated across Figs. 8, 9, 10 and 11. This control strategy not only achieves high control performance but also maintains robust effectiveness, as evidenced by the quantified variations in $C_{D}$, $C_{L}$, and the ratio of synthetic jet mass flow rate to incoming flow rate, presented in Tab. 2. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged $C_{D}$ and $C_{L}$ of the elliptical cylinder once the baseline flow has fully developed. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase. At $Ar=2$, the results of AFC control still maintain a stable baseline flow, without applying external excitation to disrupt flow stability. In the range of $Ar$ values from 1.75 to 0.75, the AFC control achieves substantial reductions in $C_{D}$ ranging from 0.9% to 15.7%, while the reduction rates in $C_{L}$ exceed 95% in all cases. Notably, the reduction in the coefficient of $C_{L}$ for elliptical cylinders, ranging from 99.7% to 95.2%, aligns with the achievement of a completely stable state (free from oscillations and fluctuations) for both $C_{D}$ and $C_{L}$ as illustrated in Fig. 10. This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig. 8.
Figure 12: Action, which is the mass flow rate of the synthetic jets. The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
@@
Interestingly, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the $C_{D}$ significantly increase to 26.9%, 35.9%, 37.6%, and 43.6% respectively, while the reduction rates of the $C_{L}$ are observed at 68.0%, 66.1%, 50.2%, and 60.0%. Protas and WesfreidProtas and Wesfreid (2002) delineated the instantaneous drag coefficient into two components: the invariant drag ($C_{D}^{b}$) and the amendable drag ($C_{D}^{0}$). The $C_{D}^{b}$ remains constant at a given Reynolds number, while the $C_{D}^{0}$ is related to oscillatory flow phenomena, representing the additional resistance caused by fluid oscillations. This amendable part can be adjusted through external interventions, demonstrating the potential for flow control. To mitigate the $C_{D}^{0}$ caused by flow instabilities, we utilize external excitations such as synthetic jets to manipulate the flow field, thereby optimizing the overall fluid dynamics performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, enhancing the controllable component of the drag. This explains why the drag reduction ratio gradually increases with the decrease in $Ar$ as shown in Tab. 2. However, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the $C_{L}$ are significantly lower than when $Ar$ is greater than or equal to 0.75. Within this range of $Ar$, only 50.2% to 68% of the $C_{L}$ is reduced, consistent with the fact that neither the $C_{L}$ nor the $C_{D}$ achieve the complete stability observed in Fig. 11. Additionally, the baseline instability is not fully mitigated as shown in Fig. 9. This highlights the fact that there is still room for improvement in the current DRL-based AFC framework for controlling the flow instability of elliptical cylinders when $Ar$ is less than or equal to 0.5. Addressing this challenge remains a significant hurdle in the field of active flow control technology for slender structures.
Fig. 12 illustrates the time history curve of the action (the mass flow rate of synthetic jets) during the application of DRL for AFC in the flow around an elliptical cylinder with $Ar$ ranging from 0 to 2. At $Ar=2$, the AFC has minimal impact, with the action value close to zero, indicating stable flow control without nonphysical perturbations. From $Ar=1.75$ to $Ar=0.75$, the agent employs a deliberate strategy of selecting larger action values for higher mass flow rates, aiming to assert flow control through significant energy inputs. As the control progresses, the action values gradually decrease, reducing the mass flow rate. In the later stages, the action values converge towards zero, optimizing for energy efficiency while maintaining flow characteristics. However, at $Ar=0.5$ and $Ar=0.25$, stable flow control is challenging, as the mass flow rate does not stabilize near zero. The non-zero action values indicate the difficulty in suppressing the chaotic wake for vertical slender bodies under these conditions. For $Ar=0.1$ and $Ar=0$, the action values exhibit periodic fluctuations, showcasing the DRL agent’s ability to recognize and exploit dominant flow frequencies.